OpenStudy (lrki):

Someone please help? (2squareroot5+3squareroot7)^2

OpenStudy (anonymous):

\[(2\sqrt5+3\sqrt7)(2\sqrt5+3\sqrt7)\] is a start then you can do the 4 multiplications first one is \(2\sqrt5\times 2\sqrt5=4\times 5=20\)

OpenStudy (whpalmer4):

It may or may not be helpful to recall that \((a+b)^2 = (a+b)(a+b) = a^2+2ab+b^2\) Then let \(a = 2\sqrt{5}\) and \(b=3\sqrt{7}\) You still have 4 multiplications, but perhaps they don't look as frightening...